OpenStudy (anonymous):

what is the sum these mixed numbers 4 1/2 + 2 1/4 =6 1/4 ? or 6 7/8

OpenStudy (turingtest):

9/2+9/4=18/4+9/4=27/4=6 3/4

OpenStudy (anonymous):

are you asking the sum of 4 and a hlaf + 2 and a quarter?

OpenStudy (turingtest):

4 1/2=9/2=18/4 2 1/4=9/4 that adds to 27/4=6 3/4 not one of the choices listed